Tom, Bill, and Brian order two large pizzas. Tom eats two slices, while Bill and Brian both have three. If two-thirds of the total amount of pizza is remaining and assuming each of the slices represents an equal portion of the pizza, how many slices were originally in each pizza?
(A) 30 (B) 24 (C) 18 (D) 12 (E) None of the above.

Answer by checkley77(12569)

IF THERE WAS 2/3 OF THE SLICES LEFT THEN THESE 3 ATE 1/3 OF THE SLICES.
2+3+3=X/3
8=X/3
X=24 TOTAL SLICES.
ANSWER B)